# GHOST: Solving the Traveling Salesman Problem on Graphs of Convex Setsthanks: Project Page: https://sites.google.com/view/ghost-gcs-tsp

### v1

@@ Abstract
We study GCS-TSP, a new variant of the Traveling Salesman Problem (TSP) defined over a Graph of Convex Sets (GCS)—a powerful representation for trajectory planning that decomposes the configuration space into convex regions connected by a sparse graph. In this setting, edge costs are not fixed but depend on the specific trajectory selected through each convex region, making classical TSP methods inapplicable. We introduce GHOST, a hierarchical framework that optimally solves the GCS-TSP by combining combinatorial tour search with convex trajectory optimization. GHOST systematically explores tours on a complete graph induced by the GCS, using a novel abstract-path-unfolding algorithm to compute admissible lower bounds that guide best-first search at both the high level (over tours) and the low level (over feasible GCS paths realizing the tour). These bounds provide strong pruning power, enabling efficient search while avoiding unnecessary convex optimization calls. We prove that GHOST guarantees optimality and present a bounded-suboptimal variant for time-critical scenarios. Experiments show that GHOST is orders-of-magnitude faster than unified mixed-integer convex programming baselines for simple cases and uniquely handles complex trajectory planning problems involving high-order continuity constraints and an incomplete GCS.
@@ Introduction
Robotic motion planning often relies on trajectory optimization, particularly in high-dimensional spaces subject to kinodynamic constraints. However, even a single obstacle in the configuration space can render the problem nonconvex and computationally challenging (ratliff2009chomp; schulman2014motion). A common approach is to reformulate such problems as convex programs for scalability. To this end, marcucci2023motion proposed the Graph of Convex Sets (GCS) framework, which decomposes the configuration space into convex regions connected via a graph structure. This representation enables transforming nonconvex trajectory planning into convex optimization subproblems while preserving feasibility and continuity.
Building on this foundation, prior work has extensively studied shortest-path problems on GCSs (marcucci2024shortest; natarajan2024ixg; tang2025space), enabling globally optimal navigation under convex constraints. Yet, many real-world robotics tasks—such as area coverage, inspection, and structured exploration—require visiting multiple regions in a cost-effective order. This naturally leads to the Traveling Salesman Problem on a GCS (GCS-TSP).
The input to GCS-TSP is a GCS whose vertices represent convex subsets of continuous configuration space and whose edges represent feasible transitions. The goal is to compute both a tour that visits each convex set at least once and a continuous trajectory that sequentially visits points in each set, subject to trajectory constraints. This problem tightly couples combinatorial tour optimization with continuous kinodynamic feasibility. Fig. 1 depicts an example solution generated by our method in a 2D environment. Applications include coverage and inspection planning, constrained multi-target visitation tasks, and task and motion planning.
Figure 1: GCS-TSP solution by GHOST, where the convex set (gray polygon) of each graph vertex must be visited at least once. The trajectory states (circles) are color-encoded (top to bottom in the right colorbar) by their timestamps.
@@
Combinatorial Optimization in GCSs: Advances in construction of collision-free convex sets (werner2024approximating; dai2024certified) have enabled efficient shortest-path solving in GCSs (marcucci2024shortest), supporting applications such as non-Euclidean motion planning (cohn2023non), quadrotor navigation (marcucci2024fast), and multi-robot spatiotemporal planning (tang2025space). marcucci2024graphs outlines a Mixed Integer Convex Program (MICP) for GCS-TSP that has limited scalability. philip2024mixed solves a TSP problem for moving targets with time windows, which is inherently restricted to Hamiltonian cycles on complete GCSs and does not involve high-order continuity or vertex repetition (or must specify repetition counts) for more general applications. Among various works on planning on GCSs, search-based ones (sundar2024graphs; chia2024gcs; bhat2025complete) typically have better efficiency and scalability. They share a hierarchical framework that separates the optimization of discrete and continuous variables, utilizing GCS convex restriction (marcucci2023motion) in the low-level for its efficiency and optimality.
@@
We prove GHOST is optimal under this search strategy. We also present a bounded-suboptimal variant that trades optimality for improved scalability, offering practical performance in large or time-constrained scenarios. We evaluate and demonstrate GHOST across various planning tasks in robotics. Compared to the unified MICP baseline, GHOST is orders of magnitude faster in simple cases and uniquely capable of solving general instances previously out of reach.
@@
Graph of Convex Sets (GCS): A GCS $G=(V,E)$ is a directed graph. Each vertex $v\in V$ is associated with a non-empty, compact convex set $\mathcal{X}_{v}\subset\mathbb{R}^{d}$. Each edge $e=(u,v)\in E$ exists iff transitions are allowed from vertex $u$ to vertex $v$. A path $\pi=(v_{1},\ldots,v_{|\pi|})$ in $G$ is a sequence of vertices such that $(v_{i},v_{i+1})\in E$ for all $i=1,\ldots,|\pi|-1$. A trajectory $\tau=(\mathbf{x}_{v_{1}},\ldots,\mathbf{x}_{v_{|\pi|}})$ conditioned on $\pi$ is a sequence of points where each $\mathbf{x}_{v_{i}}\in\mathcal{X}_{v_{i}}$. Additional trajectory constraints may be imposed to enforce kinodynamic feasibility and initial conditions. Typically, for each edge transition $(v_{i},v_{i+1})$, constraints of the form $(\mathbf{x}_{v_{i}},\mathbf{x}_{v_{i+1}})\in\mathcal{X}_{(v_{i},v_{i+1})}$ may apply, where $\mathcal{X}_{(v_{i},v_{i+1})}$ is a closed convex set. Unlike in standard graphs with fixed edge weights, the cost of traversing $\pi$ depends on the optimal trajectory conditioned on it. Specifically, a trajectory $\tau$ is defined as $c(\tau)=\sum_{i=1}^{|\pi|-1}c(\mathbf{x}_{v_{i}},\mathbf{x}_{v_{i+1}})$, where $c(\mathbf{x}_{v_{i}},\mathbf{x}_{v_{i+1}})$ is a proper, closed, convex, positive, and boundedaway-from-zero cost function. The cost of the path $\pi$ is then $c(\pi)=\min_{\tau\text{ conditioned on }\pi}c(\tau)$.
Graph problems on GCSs typically involve both discrete decisions (selecting the sets and their visitation order) and continuous decisions (choosing trajectory points within those sets). When formulated using network-flow representations (marcucci2024graphs), such problems result in mixed-integer nonconvex programs, which are generally intractable. marcucci2023motion introduced the “GCS convex restriction” trick to solve the shortest-path problem on a GCS from a given starting point $\mathbf{x}_{s}\in\mathcal{X}_{s}$ to a target point $\mathbf{x}_{t}\in\mathcal{X}_{t}$. The key idea is to fix a path $\pi$—typically a simple path (i.e., without vertex repetition)—and solved a convex program to compute the optimal trajectory conditioned on that path. Paths are explored iteratively, often guided by a convex relaxation over discrete set-selection variables. More recent GCS planners (natarajan2024ixg; chia2024gcs; tang2025space) build on the same idea but use best-first search over feasible paths, leveraging the fact that trajectory optimization along a fixed path is convex and efficient.
Traveling Salesman Problem on a GCS (GCS-TSP): GCS-TSP generalizes the classical TSP to the GCS setting. Given a GCS $G=(V,E)$, the goal is to find a tour—a closed path $\pi=(v_{1},v_{2},\ldots,v_{|\pi|})$ that visits each vertex in $\mathcal{V}$ at least once—along with its associated trajectory $\tau=(\mathbf{x}_{v_{1}},\mathbf{x}_{v_{2}},...,\mathbf{x}_{v_{|\pi|}}=\mathbf{x}_{v_{1}})$ through the convex sets that starts and ends at the same point, such that the trajectory cost is minimized. Formally, we define GCS-TSP as:
@@
$(\mathbf{x}_{v_{i-1}},\mathbf{x}_{v_{i}})\in\mathcal{X}_{(v_{i-1},v_{i})},$ $\forall i=2,\ldots,|\pi|.$ (4)
@@
In this section, we establish a static lower-bound metric for a given GCS $G=(V,E)$ to guide both levels of search in the GHOST framework. This metric is also required by the TSP subroutine (see the next section), which generates tour candidates in best-first order. The key idea is to define a lower-bound distance metric that underestimates the true trajectory cost between any vertex pair. Summing these lower bounds over an abstract path yields a provable lower bound on the cost of any abstract path. This enables GHOST to explore abstract tours in order of increasing lower-bound estimates, facilitating early pruning while preserving optimality.
@@
We construct a lower-bound graph (LBG) to efficiently estimate path costs. Unlike standard graphs where edge costs provide meaningful path estimates, the cost of traversing a GCS edge $(u,v)\in E$ depends on trajectory endpoints $\mathbf{x}_{u}\in\mathcal{X}_{u}$ and $\mathbf{x}_{v}\in\mathcal{X}_{v}$, which are unconstrained unless the full path is fixed. As a result, edge costs in $G$ yield loose lower bounds. Instead, we define the LBG on triplets $(u,v,w)$, each representing a vertex $v\in V$ along with its incoming and outgoing convex regions, allowing for a tighter lower bound that captures the “volume” of $\mathcal{X}_{v}$.
Given a GCS $G=(V,E)$, the LBG $\mathcal{H}=(\mathcal{P},\mathcal{F})$ is a directed hypergraph. A passage triplet $p=(u,v,w)\in\mathcal{P}$ is formed from every pair of adjacent edges $(u,v),(v,w)\in E$, and a hyperedge $(p,p^{\prime})\in\mathcal{F}$ connects each $p=(u,v,w)$ to $p^{\prime}=(v,w,z)$. Each triplet $p=(u,v,w)$ is labeled with a lower-bound cost $b_{p}$, computed as the optimal trajectory cost conditioned on $p$ via a convex program. Since this program relaxes any constraints on $\mathbf{x}_{u}\in\mathcal{X}_{u}$ and $\mathbf{x}_{w}\in\mathcal{X}_{w}$, $b_{p}$ never overestimates the cost portion $c(\mathbf{x}_{u},\mathbf{x}_{v})+c(\mathbf{x}_{v},\mathbf{x}_{w})$ for any trajectory conditioned on a path using passage $(u,v,w)$. Fig. 2-(a) demonstrates an example of $b_{p}$. The LBG has also been used in natarajan2024ixg to obtain an admissible heuristic for accelerating A-like search in GCSs.
@@
Any path $\pi$ on $G$ can be equivalently written as $(p_{1},\ldots,p_{k})$, where each $p_{i}=(u_{i},v_{i},w_{i})$, and $(u_{1},v_{1}=u_{2},\ldots,w_{k-1}=v_{k},w_{k})$. We define $\mathcal{L}(\pi)$ as the cost accumulating the lower-bound costs of triplets along $\pi$:
@@
Theorem 3 in natarajan2024ixg show that for any trajectory $\tau$ from $\mathcal{X}_{u}$ to $\mathcal{X}_{w^{\prime}}$ conditioned on an open path $\pi$, we have $\mathcal{L}(\pi^{*}_{u,w^{\prime}})\leq\mathcal{L}(\pi)\leq c(\tau)$, where $\pi^{*}_{u,w^{\prime}}$ is the lower-bound-cost-minimal path from $u$ to $w^{\prime}$. Theorem 1 generalizes this result to any path unfolded from an abstract path.
@@
Figure 2: LBG and path unfolding. (a) Optimal trajectory (red dashed line) establishes a lower-bound cost $b_{p}$ for the trajectories (black dashed curves) conditioned on possible passage $p$ along $\hat{\pi}$. (b)(c) Trajectories $\tau$ and $\tau^{\prime}$ (blue solid curves) conditioned on unfolded paths $\pi$ and $\pi^{\prime}$ of $\hat{\pi}$, respectively, where $\mathcal{L}(\pi)<\mathcal{L}(\pi^{\prime})$ and $c(\tau)>c(\tau^{\prime})$.
@@
3   create node $n$ with $n.\ell\leftarrow 1$, $n.\pi\leftarrow p$, and $n.g\leftarrow b_{p}$
@@
20    $n_{c}\leftarrow$ new child node of $n$ and $n_{c}.\pi\leftarrow$ append $p$ to $n.\pi$

### v2

@@ Abstract
We study GCS-TSP, a variant of the Traveling Salesman Problem (TSP) defined over a Graph of Convex Sets (GCS)—a powerful representation for trajectory planning that decomposes the configuration space into convex regions connected by a sparse graph. In GCS-TSP, edge costs are not fixed but depend on the specific trajectory passing through each convex region, making classical TSP methods inapplicable. We introduce GHOST, a hierarchical framework that optimally solves GCS-TSP by combining combinatorial tour search with convex trajectory optimization. GHOST systematically explores tours on a complete graph induced by the GCS, using a novel abstract-path-unfolding algorithm to compute admissible lower bounds that guide best-first search at both the high level (over tours) and the low level (over feasible GCS paths realizing the tour). These bounds provide strong pruning power, reducing unnecessary optimization calls. We prove that GHOST guarantees optimality and present a bounded-suboptimal variant for time-critical settings. Experiments show that GHOST is orders-of-magnitude faster than unified mixed-integer convex programming baseline while uniquely handling complex problems involving high-order continuity constraints and incomplete GCSs.
Project pagehttps://sites.google.com/view/ghost-gcs-tsp
Codehttps://github.com/reso1/ghost
@@ Introduction
Robotic motion planning often relies on trajectory optimization, particularly in high-dimensional spaces subject to kinodynamic constraints. However, even a single obstacle in the configuration space can render the problem nonconvex and computationally challenging (ratliff2009chomp; schulman2014motion). A common approach is to reformulate such problems as convex programs for scalability. To this end, marcucci2023motion proposed the Graph of Convex Sets (GCS) framework, which decomposes the configuration space into convex regions connected via a graph structure. This representation enables transforming nonconvex trajectory planning into convex optimization subproblems while preserving feasibility and continuity. Building on this foundation, prior work has extensively studied shortest-path problems on GCSs (marcucci2024shortest), enabling globally optimal navigation under convex constraints. Yet, many real-world robotics tasks—such as area coverage, inspection, and structured exploration—require visiting multiple regions in a cost-effective order. This naturally leads to the Traveling Salesman Problem on a GCS (GCS-TSP).
The input to GCS-TSP is a GCS whose vertices represent convex subsets of continuous configuration space and whose edges represent feasible transitions. The goal is to compute a tour that visits each convex set at least once and a continuous trajectory through representative points within these sets, satisfying trajectory constraints. This problem tightly couples combinatorial tour optimization with continuous kinodynamic feasibility. Fig. 1 depicts an example solution generated by our method in a 2D environment. Applications include coverage and inspection planning, constrained multi-target visitation tasks, and task and motion planning.
Figure 1: GCS-TSP solution by GHOST. Gray polygons are convex sets. Trajectory color follows the bar (top-bottom).
@@
Combinatorial Optimization in GCSs: Advances in constructing collision-free convex sets (werner2024approximating; dai2024certified) have enabled efficient shortest-path solving in GCSs (marcucci2024shortest), supporting applications such as non-Euclidean motion planning (cohn2025non), quadrotor navigation (marcucci2024fast), and multi-robot planning (tang2025space). marcucci2024graphs formulated a Mixed Integer Convex Program (MICP) for GCS-TSP, later extended to moving-target settings (philip2024mixed), but restricted to Hamiltonian cycles on complete GCSs with poor scalability. bhat2025complete proposed a two-level approach for 3D moving-target TSP, where the high level incrementally expands a partial GCS path, relying on problem-specific time-window assumptions, small branching factors, and piecewise-linear motion models.
@@
We prove GHOST is optimal under this search strategy. We also present a bounded-suboptimal variant that trades optimality for improved scalability. We evaluate and demonstrate GHOST across various planning tasks in robotics. Compared to the unified MICP baseline, GHOST is orders of magnitude faster in simple cases and uniquely capable of solving general instances previously out of reach.
@@
Graph of Convex Sets (GCS): A GCS $G=(V,E)$ is a directed graph. Each vertex $v\in V$ is associated with a non-empty, compact convex set $\mathcal{X}_{v}$. Each edge $(u,v)\in E$ exists iff transitions are allowed from vertices $u$ to $v$. A path $\pi=(v_{1},\ldots,v_{|\pi|})$ is a sequence of vertices where $(v_{i},v_{i+1})\in E$ for all $i=1,\ldots,|\pi|-1$. A trajectory $\tau=(\mathbf{x}_{v_{1}},\ldots,\mathbf{x}_{v_{|\pi|}})$ conditioned on $\pi$ is a sequence of points where each $\mathbf{x}_{v_{i}}\in\mathcal{X}_{v_{i}}$. Additional convex constraints may be imposed on the trajectory, e.g., kinodynamic feasibility. Typically, for each edge $(v_{i},v_{i+1})\in E$, constraints of the form $(\mathbf{x}_{v_{i}},\mathbf{x}_{v_{i+1}})\in\mathcal{X}_{(v_{i},v_{i+1})}$ may apply, where $\mathcal{X}_{(v_{i},v_{i+1})}$ is a closed convex set. Unlike in standard graphs with fixed edge weights, the cost of traversing $\pi$ depends on the optimal trajectory conditioned on it. Specifically, we define the trajectory cost as $c(\tau)=\sum_{i=1}^{|\pi|-1}c(\mathbf{x}_{v_{i}},\mathbf{x}_{v_{i+1}})$, where $c(\mathbf{x}_{v_{i}},\mathbf{x}_{v_{i+1}})$ is a proper, closed, convex, positive, and boundedaway-from-zero cost function.
Graph optimization problems on GCSs involve both discrete decisions (selecting sets and their visitation order) and continuous decisions (choosing trajectory points within those sets). When formulated as network flows (marcucci2024graphs), they yield mixed-integer nonconvex programs, which are generally intractable. marcucci2023motion introduced the “GCS convex restriction” trick that solves the shortest-path problem on a GCS from a given starting point $\mathbf{x}_{s}\in\mathcal{X}_{s}$ to a target point $\mathbf{x}_{t}\in\mathcal{X}_{t}$ by fixing a path $\pi$—typically simple (i.e., without vertex repetition)—and solving a convex program for the optimal trajectory conditioned on $\pi$. Paths are explored iteratively, often guided by a convex relaxation over discrete set-selection variables. Recent GCS planners (natarajan2024ixg; chia2024gcs; tang2025space) extend this idea with best-first search over feasible paths, leveraging the fact that trajectory optimization along a fixed path is convex and efficient.
Traveling Salesman Problem on a GCS (GCS-TSP): GCS-TSP generalizes the classical TSP to the GCS setting. Given a GCS $G=(V,E)$, the goal is to find a tour—a closed path $\pi=(v_{1},v_{2},\ldots,v_{|\pi|}=v_{1})$ that visits each vertex in $V$ at least once—along with its associated trajectory $\tau=(\mathbf{x}_{v_{1}},\mathbf{x}_{v_{2}},...,\mathbf{x}_{v_{|\pi|}}=\mathbf{x}_{v_{1}})$ through the convex sets that starts and ends at the same point, such that the trajectory cost is minimized. Formally, we define GCS-TSP as:
@@
$(\mathbf{x}_{v_{i}},\mathbf{x}_{v_{i+1}})\in\mathcal{X}_{(v_{i},v_{i+1})},$ $\forall i=1,\ldots,|\pi|-1.$ (4)
@@
In this section, we establish a static lower-bound metric for a given GCS $G=(V,E)$ to guide both levels of search in our GHOST framework. This metric is also required by the TSP subroutine (see next section), which generates tour candidates in best-first order. The key idea is to define a lower-bound distance metric that underestimates the true trajectory cost between any vertex pair. Summing these lower bounds over an abstract path yields a provable lower bound cost of the trajectories conditioned on it. This enables GHOST to explore abstract tours in order of increasing lower-bound estimates, enabling early pruning while preserving optimality.
@@
We construct a lower-bound graph (LBG) to efficiently estimate path costs. Unlike standard graphs where edge costs provide meaningful path estimates, the cost of traversing a GCS edge $(u,v)\in E$ depends on trajectory endpoints $\mathbf{x}_{u}\in\mathcal{X}_{u}$ and $\mathbf{x}_{v}\in\mathcal{X}_{v}$, which are unconstrained unless the full path is fixed. As a result, edge costs in $G$ yield loose lower bounds. Instead, we define the LBG on each triplet $(u,v,w)$, representing a vertex $v\in V$ along with its predecessor $u\in V$ and successor $w\in V$, allowing for a tighter lower bound that captures the “volume” of set $\mathcal{X}_{v}$.
Given a GCS $G=(V,E)$, the LBG $\mathcal{H}=(\mathcal{P},\mathcal{F})$ is a directed hypergraph. A triplet $(u,v,w)\in\mathcal{P}$ is a passage formed from every pair of adjacent edges $(u,v),(v,w)\in E$, and a hyperedge $(p,p^{\prime})\in\mathcal{F}$ connects each $p=(\cdot,u,v)$ to $p^{\prime}=(u,v,\cdot)$. Each triplet $p=(u,v,w)$ is labeled with a lower-bound cost $lb_{p}$, computed as the optimal trajectory cost conditioned on $p$ via a convex program. Since this program relaxes any constraints on $\mathbf{x}_{u}\in\mathcal{X}_{u}$ and $\mathbf{x}_{w}\in\mathcal{X}_{w}$, $lb_{p}$ never overestimates the cost portion $c(\mathbf{x}_{u},\mathbf{x}_{v})+c(\mathbf{x}_{v},\mathbf{x}_{w})$ for any trajectory conditioned on a path using triplet $p$. Fig. 2-(a) demonstrates an example of $lb_{p}$. The LBG has also been used in natarajan2024ixg to obtain an admissible heuristic for accelerating A-like search in GCSs.
Figure 2: LBG and path unfolding. (a) The optimal trajectory (red dashed line) yielding lower-bound cost $lb_{p}$ for any trajectories (black dashed curves) conditioned on each possible $p$ along $\hat{\pi}$. (b)(c) Trajectories $\tau$ and $\tau^{\prime}$ (blue solid curves) conditioned on $\pi$ and $\pi^{\prime}$ unfolded from $\hat{\pi}$, respectively.
@@
Any path $\pi$ on $G$ can be equivalently written as $(p_{1},\ldots,p_{k})$, where each $p_{i}=(u_{i},v_{i},w_{i})$, and $\pi=(u_{1},v_{1}=u_{2},w_{1}=v_{2}=u_{3},\ldots,w_{k-2}=v_{k-1}=u_{k},w_{k-1}=v_{k},w_{k})$. We define $\mathcal{L}(\pi)$ as the cost accumulating the lower-bound costs of triplets along $\pi$:
@@
Theorem 3 in natarajan2024ixg show that for any trajectory $\tau$ from $\mathcal{X}_{u}$ to $\mathcal{X}_{w^{\prime}}$ conditioned on an open path $\pi$, we have $\mathcal{L}(\pi^{*}_{u,w^{\prime}})\leq\mathcal{L}(\pi)\leq c(\tau)$, where $\pi^{*}_{u,w^{\prime}}$ is the lower-bound-cost-minimal path from $\mathcal{X}_{u}$ to $\mathcal{X}_{w^{\prime}}$. Theorem 1 generalizes this result to any path unfolded from an abstract path.
@@
Path Unfolding
We present our (abstract-)path-unfolding algorithm, which uses a multi-label A-like best-first search (grenouilleau2019multi) to enumerate all paths unfolded from a given abstract path $\hat{\pi}=(v_{1},v_{2},\ldots,v_{k})$ on $\hat{G}$ in non-decreasing order of their lower-bound costs. Each search state consists of a triplet $p$ and a label $\ell$, indicating progress toward the next endpoint $v_{\ell+1}$. To allow for cyclic and next-best paths to a state, each search node $n$ consists of its path $n.\pi$ (instead of a triplet) and label $n.\ell$. The node also stores the current cost $n.g$ and an estimated final cost $n.f=n.g+h_{\hat{\pi}}(p,n.\ell)$, where $h_{\hat{\pi}}(p,n.\ell)$ is an admissible heuristic underestimating the cost to sequentially visit all remaining endpoints $v_{\ell+1},\ldots,v_{k}$. Fig. 2-(bc) shows an example in which the optimal trajectory conditioned on $\pi$ can have a smaller cost than $\pi^{\prime}$, even when $\mathcal{L}(\pi)<\mathcal{L}(\pi^{\prime})$. This motivates exhaustively exploring all unfolded paths for $\hat{\pi}$ to guarantee finding the one that yields the optimal trajectory.
Pseudocode (Alg. 1) The search maintains an OPEN list of nodes $n$, prioritized by $n.f$ (Line 1). Nodes are initialized for every triplet $p$ that matches the prefix of $\hat{\pi}$, with their labels set according to how many initial vertices of $\hat{\pi}$ are matched (Lines 1-1). At each step, $n$ with the lowest $n.f$ is popped (Line 1). For an open $\hat{\pi}$, if $n.\pi$ ends at $v_{k}$ and $n.\ell=k$, the search outputs $n.\pi$ as the next-best unfolded path (Line 1). For a closed $\hat{\pi}$, if $n.\pi$ starts and ends at the same triplet $(\cdot,v_{k},\cdot)$ and $n.\ell=k$, the search outputs $n.\pi$ as the next-best unfolded path (Line 1), with $n.g$ correctly capturing the wraparound triplet cost. The search then branches on all triplets $p$ consistent with the suffix of $n.\pi$. Fig. 3 shows an example. Each child node $n_{c}$ increments the label if appending $p$ reaches endpoint $v_{n.\ell+1}$ (Line 1), updates $n_{c}.g$ by adding $lb_{p}$, and sets $n_{c}.f$ as the sum of $g$ and an admissible heuristic value $h_{\hat{\pi}}$ (Line 1). Optionally, nodes with $n_{c}.f>\bar{c}$ are pruned (Line 1), supporting efficient hierarchical search in GHOST. We employ a multi-label heuristic inspired by li2021lifelong; zhong2022optimal as defined in Lemma 2. Fig. 3 illustrates Alg. 1 with a concrete example.
Lemma 2.
Proof 2.
Consider any open path $\pi=\pi_{v_{1},v_{2}}\circ\ldots\circ\pi_{v_{k-1},v_{k}}$ unfolded from $\hat{\pi}$, where each subpath $\pi_{v_{i},v_{i+1}}=(v_{i}=w^{(i)}_{1},w^{(i)}_{2},\ldots,w^{(i)}_{\mathcal{K}_{i}}=v_{i+1})$. Now, assume the triplet $p=(u,v,w)$ at label $n.\ell$ is followed by some $p^{\prime}=(v,w,\cdot)$. The cost-to-go from $p$ can be decomposed as $lb_{p^{\prime}}+\mathcal{L}(\pi_{w,v_{\ell+1}})+\sum_{i=\ell+1}^{k-1}\left(lb_{(w^{(i)}_{\mathcal{K}_{i}-1},v_{i},w^{(i+1)}_{2})}+\mathcal{L}(\pi_{v_{i},v_{i+1}})\right)$, which is at least $\mathcal{L}(\pi^{*}_{w,v_{\ell+1}})+\sum_{i=\ell+1}^{k-1}\mathcal{L}(\pi^{*}_{v_{i},v_{i+1}})=h_{\hat{\pi}}(p,n.\ell)$. For closed paths, the inequality still holds with an additional non-negative term $lb_{(w^{(k-1)}_{\mathcal{K}_{k-1}-1},v_{k},w^{(1)}_{2})}$ on the left. In summary, $h_{\hat{\pi}}$ aggregates the minimum possible subpath costs but does not account for the triplet costs for $p^{\prime}$ and trajectory endpoints in the actual cost-to-go.
Figure 3: Illustration of Alg. 1 for $\hat{\pi}=(v_{1},\ldots)$ on the GCS from Fig. 2: OPEN evolution (nodes in gray blocks sorted top-bottom in non-decreasing $f$), node expansions (dashed arrows), and appended triplets (blue).
@@
3   create node $n$ with $n.\ell\leftarrow 1$, $n.\pi\leftarrow p$, and $n.g\leftarrow lb_{p}$
@@
20    $n_{c}\leftarrow$ new child node of $n$ with $n_{c}.\pi$ appending $p$ to $n.\pi$